OpenStudy (anonymous):

what is the equation of the normal to the curve y=x^2 at (2,3)

OpenStudy (anonymous):

\[ y= x^2\\ y'= 2 x\\ y'(2)=4 \] What is the slope of the normal?

OpenStudy (anonymous):

4 is the slope of the tangent

OpenStudy (anonymous):

-1/4 is the slope of the normal

OpenStudy (anonymous):

(2,3) \[ y-3 =-\frac 1 4(x-2) \] is the equation of the normal

OpenStudy (anonymous):

\[ y =-\frac x 4+ \frac 12 +3\\ y =-\frac x 4+ \frac 72 \\ \]

OpenStudy (anonymous):

Did you understand it?

OpenStudy (anonymous):

seems @eliassaab provided the whole solution ::-P